OpenStudy (anonymous):

Trigonometric substitution problem \[\int{\frac{{{x}^{2}}}{\sqrt{16-{{x}^{2}}}}}\]

OpenStudy (anonymous):

try either \(x=2\sin(\theta)\) or \(u=2\cos(\theta)\) either should work

OpenStudy (anonymous):

ok that was silly, i meant \[u=4\sin(\theta)\] because then \(\sqrt{16-(4\sin(\theta))^2}=\sqrt{16-16\sin^2(\theta)}=4\cos(\theta)\)

OpenStudy (anonymous):

why a 2 upfront whats the logic ?

OpenStudy (anonymous):

a mistake on my part, not logical at all

OpenStudy (anonymous):

I tried 4sin(theta) but nothing at all

OpenStudy (anonymous):

i wrote it out above

OpenStudy (anonymous):

4*INT(sinu/cosu )

OpenStudy (anonymous):

thats basically where I stuck forever

OpenStudy (anonymous):

you should have \[16\int\sin^2(\theta)d\theta\]

OpenStudy (anonymous):

true

OpenStudy (anonymous):

don't forget if \(u=4\sin(\theta)\) then \(du=4\cos(\theta)d\theta\) and the cosines cancel top and bottom

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

sorry for bothering you again @satellite73 but can you tell me the next 1-2 steps ? :D